OpenStudy (anonymous):

I have no idea where to start on this problem..... Any help, please :) -10 - 3/7 x = 5

OpenStudy (dumbcow):

i am you need to solve for x

OpenStudy (anonymous):

multiply each side of the equation with 7x

OpenStudy (anonymous):

do you need to go back to school for english, dumbcow?

OpenStudy (anonymous):

so how do I start that? do I add 3/7 to the right side?

OpenStudy (anonymous):

nurse, is x in the denominator of the fraction?

OpenStudy (anonymous):

why multiply 7?

OpenStudy (dumbcow):

yes thats why i do math

OpenStudy (anonymous):

get ride of the fractions and multiply by 7 in each side

OpenStudy (anonymous):

You get -10-3=5(7x)

OpenStudy (anonymous):

no X is not the denominator

OpenStudy (anonymous):

The fraction is 3/7

OpenStudy (anonymous):

Frederick, what do I do next?

OpenStudy (dumbcow):

-10 -(3/7)*x = 10 Add 10 to both sides _>-(3/7)*x = 20 multiply by reciprocal of -3/7 on both sides ->x = 20*(-7/3) -> x = -140/3

OpenStudy (anonymous):

you would not get -10-3=5(7x), you would get -10(7x)-3=5(7x), which does not help. Isolate x, => -(3/7)x=5-10 => x = (5-10)/(-3/7)

OpenStudy (anonymous):

nurse just google it since dumbcow doesn't know what he's doing

OpenStudy (anonymous):

sorry i made a mistake. the equation is -10-(3/7)x = 5, add 10 to both sides, then divide by -(3/7)

OpenStudy (dumbcow):

haha oops yeah i wrote the problem down wrong x=-105/3

OpenStudy (anonymous):

x=35

OpenStudy (gw2011):

Add +10 to both sides of the equation because you want to keep the unknowns (x in this case) on one side of the equation and the knowns on the otherside of the equation. This will give you the following: -3/7x=5+10=15 Now multiply both side of the equation by -7 (since there is a minus sign and you want to put the knowns on the same side of the equation) and divde by (for the same reason of putting the knowns on the same side of the equatioin). This gives you the following: x=[(15) (-7)]/3 x=-35

OpenStudy (gw2011):

I inadvertently left out the number 3 when I said to "and divide by (for the same reason...)." Obviously, the number 3 should have been placed after "divide by." However, you can see this in the equation where the number 3 is dividing the equation.